On a molar basis, air is about 78% N2 . What is the Henry's law constant for , N2 , in water if the solubility of, N2 , in water is .000535 mole/ L, on a day when the atmospheric pressure is 745 torr?
Answer #1

First convert pressure into atm

P=745/760

=.98

S=k*p

Solubility=Henry constant*pressure

kh=S/p

=.000535/.98

=5.45*10-4

This is the value of Henry constant